OpenStudy (ohmyquaffle):

Two researchers are studying the decline of orangutan populations. In one study, a population of 784 orangutans is expected to decrease at a rate of 25 orangutans per year. In a second study, the population of a group of 817 orangutans is expected to decrease at a rate of 36 per year. After how many years will the two populations be the same? The two populations will be the same after ___ years.

OpenStudy (evoker):

basically you need two linear equation to describe each the one would be y=-25x+784 and the other would be y=-36x+817. Then to find when they will have the same y we set them equal to each other or -25x+784=-36x+817 rearranging yields11x=33 or after x=3 years they will have the same pop.

OpenStudy (ohmyquaffle):

thank you so so much!

OpenStudy (welshfella):

equation relating population and years for the first one is y = 784 - 25x can you figure out what the equation for the second lot?

OpenStudy (welshfella):

@Evoker please dont do the question for the poster. Try to get him involved in the solution. he/ she will learn more that way.

OpenStudy (evoker):

I personally believe students should have multiple questions on a topic, as long as explanations are given in detail, at least one fully explained example is useful for students.